Circular No. 8751

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

(120347) 2004 SB_{60}

K. S. Noll, Space Telescope Science Institute; H. F. Levison, Southwest Research Institute; D. C. Stephens, Johns Hopkins University; and W. M. Grundy, Lowell Observatory, report the detection of a binary companion to the transneptunian object (120347) 2004 SB_{60} (MPEC 2005-J26, 2005-J28, MPO 88617; orbital elements a = 41.97 AU, e = 0.109, i = 23^{o}.92, epoch 2006 Sept. 22.0 TT). The observations were made during 2006 July 21.900– 21.925 UT with the High Resolution Camera of the Advanced Camera for Surveys on the Hubble Space Telescope (HST), using the clear filters with one 300-s exposure at each of four dithered positions on the detector. The two components are clearly resolved in each image and in the coadded image, separated by an angular distance of 0′′.110 ± 0′′.002 and differing in brightness by 2.3 magnitudes. The fainter component lies at a position angle of 154^{o}.8 ± 0^{o}.6 from the primary. The projected separation of the objects in the sky plane is 3440 ± 60 km. The HST corrected for parallax and tracked both components of (120347) as they moved together at an average rate of 0′′.034/min.

SUPERNOVAE 2006gi AND 2006gj Two apparent supernovae have been discovered on unfiltered CCD im-

ages: 2006gi by K. Itagaki (cf. IAUC 8744; communicated by S. Nakano, Sumoto, Japan), and 2006gj by T. Puckett and D. Toth (cf. IAUC 8750).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006gi Sept. 18.78 10^{h}16^{m}46.^{s}76 +73^{o}26′26′′.4 16.3 30′′ W, 144′′ N
2006gj Sept. 18.38 3 17 35.80 − 1 41 30.2 17.8 19′′.3 E, 12′′.1 N

Additional magnitudes from Itagaki for 2006gi in NGC 3147: Mar. 22, [19.0; June 3, [19.0; Sept. 19.399 UT, 16.4. Nothing is visible at the location of 2006gi on the Digitized Sky Survey. Additional magnitudes for 2006gj in UGC 2650: 2005 Oct. 21, [19.3 (Puckett); 2006 Sept. 19.13, 17.7 (D. Toth, 0.35-m TIE telescope, Las Campanas Observatory).

V733 CEPHEI N. N. Samus, Institute of Astronomy, Russian Academy of Sciences,

reports that the variable star reported on IAUC 8441 by Persson et al. has been given the designation V733 Cep.

2006 September 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green